wall paper,312 cm long and 25 cm wide is required to cover the walls of a room. length of the room is 7 m and its breadth is twice height.determine the height of the room.
1
Expert's answer
2013-01-08T08:59:21-0500
We assume, that room doesn't have any windows and doors.
For paper S=312*25
For walls S=2*(l+b)*h
l=700 sm
b=2h
2*(700+2h)h=312*25
4h^2+700h-7800=0
h^2+125h-1800=0
Hence, h=13 sm
So, we see, that input data are not right.
If we will take wall paper 3120 sm length, we get h=185 sm, that is normal result.